PHY 201 Lab Manual (Spring 2019 version) For this page: The velocity (yes, velocity, not speed) of the person increases at a constant rate (ie, linearly) from-2.0 m/s to +2.0 m/s during the time interval from t = 0.0 s to t = 1.0 s. 33. What is the acceleration of the object? Show work B Write out explicitly the function r(t). . Draw the graph of the function v(t). Include units, labels, and numerical values on the graph D. Draw the graph of the function speed(). Include units, labels, and numerical values on the graph PHY 201 Lab Manual
Answer #1

Answer :-

The velocity of person increases linearly with time,

So v2=2m/s, t2=1 sec

V1=-2m/s, t1=0sec

So answer of Q.No 33 is,

a=(v2-v1)/(t2-t1)

a=(2-(-2))/(1-0)

So a=4m/s2.

for question。 93.B z-지 at 1 0 20 -S -20 rquestion o33.d ase@apeedーガme grapt>zditわcoastaot 8peed. oe keoco at speed at eertain time is ומ tudeaf velocity at that time the speed is oo notant-→ the 8.peed.-ti tine(モ) fig : 8peed-troe graph for const speedCase (2) Speed-time graph witわconst accelern color accelera Hoの is aoastaor, A the roittal speed of Particle s eov , so speed oF Speed 0 glope= acci- time (t) case epeed-time graph cincressing ace speed tine Csec